OpenStudy (somegirl1999):

Find the length of the missing side of the right triangle. (1 point) 289 cm 17 cm 23 cm 4.79 cm

OpenStudy (somegirl1999):



OpenStudy (raden):

use the pythagorean theorem : x^2 = 8^2 + 15^2 solve for x

OpenStudy (raden):

i assumed that the length of hypotenuse is x

OpenStudy (somegirl1999):

thanx guys :)

OpenStudy (raden):

x^2 = 8^2 + 15^2 x^2 = 64 + 225 x^2 = 289 x = sqrt(289) = 17

OpenStudy (somegirl1999):

Thank you

OpenStudy (raden):

you're welcome